# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ A.5 Tree-Based wavelet
As shown in [Gavish et al., 2010], Haar wavelets can be constructed directly and adaptively from trees in a straightforward manner. Without loss of generality, we focus here on the Haar basis induced by the feature tree, where the
@@
resulting wavelet filter is subsequently applied to the samples. Symmetrically, one can construct a Haar basis from the sample tree and apply the corresponding wavelet filter to the features.
@@
where $Q^{\ell}$ is the orthogonal complement of $S^{\ell}$. As the tree $\mathcal{T}(\mathbf{M}_{c})$ is a full binary tree, the Haar-like basis constructed from the tree is essentially the standard Haar basis [Haar, 1911], denoted by $\{\bm{\beta}_{\ell,\varepsilon}\in\mathbb{R}^{m}\}$. At level $\ell$, a subtree $\widehat{\Upsilon}_{c}(\ell,s)$ splits into two sub-subtrees $\widehat{\Upsilon}_{c}(\ell+1,s_{1})$ and $\widehat{\Upsilon}_{c}(\ell+1,s_{2})$. A zero-mean Haar wavelet $\bm{\beta}_{\ell,s}\in\mathbb{R}^{m}$ has non-zero values only at the indices corresponding to leaves in the sub-subtrees and is piecewise constant on each of them. The set of these Haar wavelets, along with a constant vector $\bm{\beta}_{0}$, is complete and forms an orthonormal Haar basis. We collect all of these basis vectors as columns of matrix, denoted by $\mathbf{B}\in\mathbb{R}^{m\times m}$. Fig. 3 illustrates a Haar basis induced by a binary tree.
Proposition A.1 (Function Smoothness and Coefficient Decay [Gavish et al., 2010]).
@@
Prop. A.2 implies that with $L_{1}$ approximation, it is sufficient to estimate the coefficients for function approximation. In the Haar domain, estimating these coefficients can be achieved using the fast wavelet transform [Beylkin et al., 1991].
@@ A.6 Wasserstein distance
The Wasserstein distance [Villani, 2009] measures the discrepancy between two probability distributions over a given metric space. It is related to optimal transport (OT) theory [Monge, 1781, Kantorovich, 1942], where the goal is to quantify the cost of transforming one distribution into another.
@@
for all measurable sets $A,B\subseteq\mathcal{X}$. It represents the minimal “cost” to transport the mass of $\mu$ to $\nu$, which is also known as “earth mover’s distance” [Peyré et al., 2019, Gavish et al., 2010].
@@
When computing the Wasserstein distance between discrete probability distributions, the computational bottleneck often lies in constructing and processing the ground pairwise distance matrix. This matrix encodes the distances between every pair of discrete support points in the distributions, requiring $O(m^{2})$ storage and $O(m^{3}\log m)$ computation [Peyré et al., 2019]. This complexity arises due to solving a linear program for the optimal transport problem, which scales poorly with the number of points $m$. Consequently, applying optimal transport directly becomes infeasible for large-scale datasets [Bonneel et al., 2011].
Several approximations have been proposed to reduce this computational complexity. The Sinkhorn distance [Cuturi, 2013, Chizat et al., 2020] introduces an entropy regularization term to the objective function, enabling the use of iterative matrix scaling algorithms in quadratic. Graph-based methods exploit sparsity in the transport graph to simplify the problem structure, while sampling-based approaches approximate distributions using a subset of support points, reducing computational demand [Gasteiger et al., 2021, Sommerfeld et al., 2019, Altschuler et al., 2017].
@@ A.7 Tree-Wasserstein distance
The Wasserstein distance requires solving a linear programme whose computational cost is quadratic in the number of support points. A widely used strategy to mitigate this computation cost is to approximate the ground metric with a tree metric, thereby defining the tree-Wasserstein distance (TWD) [Indyk and Thaper, 2003]. Because transport on a tree admits a closed-form solution, TWD can be computed in $O(m)$ time for a tree with $m$ leaves, which makes it attractive for large-scale applications.
Let $T=(V,E,\mathbf{A})$ be a tree with $N_{\text{leaf}}$ leaves. The tree distance $d_{T}:V\times V\rightarrow\mathbb{R}$ is the sum of weights of the edges on the shortest path between any two nodes on $T$. Let $\Upsilon_{T}(v)$ be the set of nodes in the subtree of the tree $T$ rooted at the node $v\in V$. For any node $u\in V$, there exists a unique parent $v$ s.t. the weight is defined by the tree distance, i.e., $w_{u}=d_{T}(u,v)$ [Bondy and Murty, 2008]. Given two discrete distributions $\bm{\rho}_{1},\bm{\rho}_{2}\in\mathbb{R}^{N_{\text{leaf}}}$ supported on the tree $T$, the TWD is formally defined by
$\mathrm{TW}(\bm{\rho}_{1},\bm{\rho}_{2},T)=\textstyle\sum_{v\in V}w_{v}\left\lvert\sum_{u\in\Upsilon_{T}(v)}\left(\bm{\rho}_{1}(u)-\bm{\rho}_{2}(u)\right)\right\rvert,$ (30)
@@
It is shown in [Le et al., 2019, Yamada et al., 2022] that for a tree $T$ and probability measures supported on the tree $T$, the TWD computing using $T$ is the Wasserstein distance on the tree metric $d_{T}$, i.e.,
@@
However, it is important to note that most of the TWD methods [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019, Chen et al., 2024] are designed to approximate a Wasserstein distance with an Euclidean ground metric. Specifically, their goal is
@@ A.8 Graph convolutional networks
Graph Convolutional Networks (GCNs) [Kipf and Welling, 2016] have gained significant attention in graph machine learning, where nodes in a graph are typically assumed in Euclidean spaces. By generalizing convolutional operations to graphs, GCNs effectively capture the dependencies between nodes. This capability enables high accuracy in tasks such as node classification, link prediction, and graph classification, making GCNs useful in applications like social networks [Hamilton et al., 2017], molecular structures [Duvenaud et al., 2015, Gilmer et al., 2017], knowledge graphs [Wang et al., 2019], recommendation systems [Ying et al., 2018], and drug discovery [Wu et al., 2018]. A brief overview of the GCN framework is presented below.
@@
$\mathbf{h}_{i}^{(\ell)}=\sigma\left(\mathbf{h}^{(\ell)}_{i}+\sum_{s\in[[i]]}w_{is}\mathbf{h}^{(\ell)}_{s}\right),$ (34)
@@ A.9 Hyperbolic graph convolutional networks
Hyperbolic Graph Convolutional Networks (HGCNs) [Chami et al., 2019, Dai et al., 2021] generalize Graph Convolutional Networks (GCNs) [Kipf and Welling, 2016] to hyperbolic spaces, using the inductive bias of hyperbolic geometry to better capture hierarchical structures in graph data. By embedding nodes in a negatively curved space, HGCNs provide improved representational capacity for graphs with hierarchies, while preserving the scalability of standard GCNs. This extension requires adapting the steps in GCNs, including feature transformation and neighborhood aggregation, to conform with the geometry of hyperbolic space. HGCNs have shown empirical advantages across a range of tasks, including social network analysis, recommendation systems, and biological network modeling, demonstrating the effectiveness of geometric inductive biases in deep learning on non-Euclidean domains. We briefly review the HGCNs framework below.
@@
Mapping Euclidean node features to hyperbolic representations. Following [Chami et al., 2019], HGCNs begin by mapping input Euclidean node features to a Lorentz model using the exponential map at the origin. Specifically, given a Euclidean feature vector $\mathbf{x}_{i}\in\mathbb{R}^{m}$, its hyperbolic embedding is initialized as
@@
Hyperbolic feature transformation. To perform feature transformations in hyperbolic space, HGCNs [Chami et al., 2019] use the logarithmic and exponential maps to move between the manifold and the tangent space at the origin. Let $\mathbf{W}\in\mathbb{R}^{m^{\prime}\times m}$ be a learnable weight matrix. The hyperbolic equivalent of matrix multiplication is defined as
@@
Hyperbolic neighborhood aggregation. Neighborhood aggregation in HGCNs is performed in the tangent space and leverages hyperbolic attention to account for the hierarchical relationships within the graph. Given hyperbolic embeddings $\mathbf{x}_{i}^{H}$ and $\mathbf{x}_{j}^{H}$, both are first mapped to the tangent space at the origin. The attention weights are computed via an Euclidean Multi-Layer Perceptron ($\mathrm{MLP}$) applied to the concatenation of the tangent vectors
@@
where $[[i]]$ denotes the neighbors of node $i$, and ||||| | indicates concatenation. Aggregation is then performed via a hyperbolic weighted average in the tangent space of $\mathbf{x}_{i}^{H}$
@@
Remark.
In the hyperbolic feature transformation step of HGCNs [Chami et al., 2019], the standard approach involves mapping points between the hyperbolic manifold and the Euclidean tangent space via logarithmic and exponential maps. This allows feature transformations to be performed using Euclidean operations, such as linear transformations followed by bias addition. However, this design only partially uses the advantage of the geometry of hyperbolic space, as the actual transformation takes place in the flat tangent space rather than directly on the curved manifold. To exploit the structure of hyperbolic space, [Dai et al., 2021] proposed a fully hyperbolic feature transformation. Specifically, they introduce a Lorentz linear transformation that operates directly in the hyperbolic space
@@
Appendix B Additional related works
We discuss additional related works and their objectives.
@@ Appendix B Additional related works
Manifold learning with Wasserstein distance. Manifold learning is a nonlinear dimensionality reduction approach designed for high-dimensional data that intrinsically lie on a lower-dimensional manifold [Fefferman et al., 2016, Coifman and Lafon, 2006, Belkin and Niyogi, 2008]. Most classical methods derive manifold representations from discrete data samples by constructing a graph, where nodes correspond to data points and edges reflect pairwise similarities. These similarities are often defined using a Gaussian affinity kernel based on the Euclidean norm between the data samples. Recent studies have demonstrated the benefits of using the Wasserstein distance as a more geometry-aware alternative to Euclidean distances in manifold learning frameworks [Gavish et al., 2010, Ankenman, 2014], where the Wasserstein distance incorporates the feature relationship. This direction has shown promising results in applications such as matrix organization [Mishne et al., 2017], analysis of neuronal activity patterns [Mishne et al., 2016], and molecular shape space modeling [Kileel et al., 2021]. Building on these ideas, our work explores the integration of the diffusion operator [Coifman and Lafon, 2006] with the TWD to jointly learn hierarchical representations for both samples and features.
@@
Tree-Wasserstein distance. The Tree-Wasserstein Distance (TWD) offers a computationally efficient alternative to the Wasserstein distance by approximating the ground cost metric with a tree structure. It is designed to compare probability distributions by quantifying the amount of mass that must be transported between them, where the transport is governed by distances along a tree. By approximating the original Euclidean ground metric with a tree metric, TWD enables faster computation while maintaining a close approximation to the true Wasserstein distance. This trade-off between efficiency and fidelity has been validated in a range of studies [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019], where TWD has been shown to provide a reliable and scalable surrogate for optimal transport in Euclidean settings. For instance, the Quadtree [Indyk and Thaper, 2003] recursively partitions the ambient space into hypercubes to build random trees, which are then used to calculate the TWD. Flowtree [Backurs et al., 2020] refines this approach by focusing on optimal flow and its cost within the ground metric. Similarly, the Tree-Sliced Wasserstein Distance (TSWD) [Le et al., 2019] further improves robustness by averaging TWD values computed over multiple randomly sampled trees, with variants such as TSWD-1, TSWD-5, and TSWD-10 corresponding to the number of sampled trees used. Recent advancements include WQTWD and WCTWD [Yamada et al., 2022], which employ Quadtree or clustering-based tree structures and optimize tree weights to approximate the Wasserstein distance. UltraTree [Chen et al., 2024] introduces an ultrametric tree by minimizing OT regression cost, aiming to approximate the Wasserstein distance while respecting the original metric space.
@@
Most recently, [Lin et al., 2025] proposed a tree construction method in TWD literature that differs fundamentally from existing TWD approaches [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019]. Whereas conventional TWD methods focus on how TWD can be close to the true Wasserstein distance with the Euclidean ground metric, the method in [Lin et al., 2025] aims to efficiently compute a Wasserstein distance with a ground metric that recovers a latent hierarchical structure underlying hierarchical high-dimensional features. Unlike prior TWD methods that treat tree construction as a heuristic or approximation step to Euclidean distances, the approach in [Lin et al., 2025] treats the tree as a geometric object that reflects intrinsic hierarchical relationships among features. In particular, the constructed tree is not merely a proxy for an ambient Euclidean metric, but rather a data-driven structure whose tree distances approximate geodesic paths on a latent, unobserved hierarchical metric space. This allows the resulting TWD to encode meaningful relationships in settings where the data are governed by hidden hierarchies. In our work, we adopt the tree construction method from [Lin et al., 2025], as it provides a principled and geometry-aware foundation for learning hierarchical representations of both samples and features.
@@
Sliced-Wasserstein distance. Another widely used approach for efficiently approximating the Wasserstein distance is the Sliced-Wasserstein Distance (SWD) [Rabin et al., 2012, Bonneel et al., 2015]. Instead of solving the high-dimensional optimal transport problem directly, SWD projects the input distributions onto one-dimensional subspaces along random directions, where the Wasserstein distance admits a closed-form solution. By averaging the results over many such slices, it provides a meaningful measure of similarity while retaining much of the geometric structure inherent in the distributions. This approach is advantageous in high-dimensional settings, where computational efficiency is critical, making it popular in applications such as generative modeling [Kolouri et al., 2019], domain adaptation [Bonet et al., 2023b], and statistical inference [Bonet et al., 2023a].
@@
Co-Manifold learning. Co-manifold learning aims to simultaneously uncover the manifolds of both samples and features in a data matrix by treating their relationships as mutually informative. That is, the geometry of the features is informed by the samples, and vice versa. This joint modeling approach has been widely explored in tasks such as joint embedding [Ankenman, 2014, Gavish and Coifman, 2012, Yair et al., 2017, Leeb and Coifman, 2016] and dimensionality reduction [Mishne et al., 2016, Shahid et al., 2016, Mishne et al., 2019], under the assumption that both samples and features lie on smooth, low-dimensional manifolds. By iteratively refining representations across both modes, co-manifold methods aim to capture the manifolds of both samples and features. These approaches typically assume that the manifolds involved are Riemannian and characterized by non-negative curvature [Leeb and Coifman, 2016, Fefferman et al., 2016]. However, such assumptions may be limiting in scenarios where the data exhibit hierarchical organization, which is more naturally modeled in negatively curved spaces such as hyperbolic geometry.
@@
Unsupervised ground metric learning. Ground metric learning [Cuturi and Avis, 2014] focuses on learning a distance function that defines the cost matrix in the optimal transport problem, thereby capturing meaningful relationships between data elements. The choice of ground metric plays a critical role in downstream tasks [Heitz et al., 2021, Zen et al., 2014] such as clustering, transport-based inference, and dimensionality reduction. Many existing approaches rely on prior knowledge of the cost matrix structure or access to label information to guide the learning process [Wang and Guibas, 2012, Li et al., 2019, Stuart and Wolfram, 2020]. In many cases, however, labeled data is unavailable, making it necessary to develop unsupervised methods that rely solely on the data itself.
Unsupervised ground metric learning seeks to infer a distance metric that captures meaningful pairwise relationships directly from the data, without relying on external labels or prior knowledge. A recent work in this direction is the Wasserstein Singular Vectors (WSV) method introduced by [Huizing et al., 2022], which jointly computes Wasserstein distances over samples and features by using the Wasserstein distance matrix in one domain (e.g., samples) as the ground metric for computing distances in the other (e.g., features). However, the method is computationally demanding due to the repeated evaluation of high-dimensional Wasserstein distances. To address this issue, the recent work [Düsterwald et al., 2025] proposed Tree-WSV, which uses tree-based approximations of the Wasserstein distance to reduce computational cost. Both WSV and Tree-WSV adopt an alternating scheme that iteratively updates distances across rows and columns, a strategy that aligns conceptually with our framework.
@@
In contrast, our method is explicitly designed to represent the hierarchical structures for both rows and columns. The key distinction lies in our use of the tree in the TWD following [Lin et al., 2025]: we construct the tree using the method of [Lin et al., 2025], which employs hyperbolic embeddings and diffusion densities to reflect meaningful hierarchical relationships among data points. We then integrate this specific tree-based TWD into the construction of a diffusion operator, whose diffusion densities are used to generate hyperbolic embeddings that encode the joint structure of both samples and features. In doing so, we not only model each hierarchical structure but also capture their mutual influence through an alternating learning process.

### v3

@@ A.5 Tree-Based wavelet
Haar wavelets can be constructed directly and adaptively from trees in a straightforward manner [43]. Without loss of generality, we focus here on the Haar basis induced by the feature tree, where the resulting wavelet filter is subsequently applied to the samples. Symmetrically, one can construct a Haar basis from the sample tree and apply the corresponding wavelet filter to the features.
@@
where $Q^{\ell}$ is the orthogonal complement of $S^{\ell}$. As the tree $\mathcal{T}(\mathbf{M}_{c})$ is a full binary tree, the Haar-like basis constructed from the tree is essentially the standard Haar basis [40], denoted by $\{\bm{\beta}_{\ell,\varepsilon}\in\mathbb{R}^{m}\}$. At level $\ell$, a subtree $\widehat{\Upsilon}_{c}(\ell,s)$ splits into two sub-subtrees $\widehat{\Upsilon}_{c}(\ell+1,s_{1})$ and $\widehat{\Upsilon}_{c}(\ell+1,s_{2})$. A zero-mean Haar wavelet $\bm{\beta}_{\ell,s}\in\mathbb{R}^{m}$ has non-zero values only at the indices corresponding to leaves in the sub-subtrees and is piecewise constant on each of them. The set of these Haar wavelets, along with a constant vector $\bm{\beta}_{0}$, is complete and forms an orthonormal Haar basis. We collect all of these basis vectors as columns of a matrix, denoted by $\mathbf{B}\in\mathbb{R}^{m\times m}$. Fig. 3 illustrates a Haar basis induced by a binary tree.
Proposition A.1 (Function Smoothness and Coefficient Decay [43]).
@@
Prop. A.2 implies that with $L_{1}$ approximation, it is sufficient to estimate the coefficients for function approximation. In the Haar domain, estimating these coefficients can be achieved using the fast wavelet transform [117].
@@ A.6 Wasserstein distance
The Wasserstein distance [46] measures the discrepancy between two probability distributions over a given metric space. It is related to optimal transport (OT) theory [47, 48], where the goal is to quantify the cost of transforming one distribution into another.
@@
for all measurable sets $A,B\subseteq\mathcal{X}$. It represents the minimal “cost” to transport the mass of $\mu$ to $\nu$, which is also known as “earth mover’s distance” [49, 43].
@@
When computing the Wasserstein distance between discrete probability distributions, the computational bottleneck often lies in constructing and processing the ground pairwise distance matrix. This matrix encodes the distances between every pair of discrete support points in the distributions, requiring $O(m^{2})$ storage and $O(m^{3}\log m)$ computation [49]. This complexity arises due to solving a linear program for the optimal transport problem, which scales poorly with the number of points $m$. Consequently, applying optimal transport directly becomes infeasible for large-scale datasets [118].
Several approximations have been proposed to reduce this computational complexity. The Sinkhorn distance [119, 120] introduces an entropy regularization term to the objective function, enabling the use of iterative matrix scaling algorithms in quadratic. Graph-based methods exploit sparsity in the transport graph to simplify the problem structure, while sampling-based approaches approximate distributions using a subset of support points, reducing computational demand [121, 122, 123].
@@ A.7 Tree-Wasserstein distance
The Wasserstein distance requires solving a linear programme whose computational cost is quadratic in the number of support points. A widely used strategy to mitigate this computation cost is to approximate the ground metric with a tree metric, thereby defining the tree-Wasserstein distance (TWD) [32]. Because transport on a tree admits a closed-form solution, TWD can be computed in $O(m)$ time for a tree with $m$ leaves, which makes it attractive for large-scale applications.
Let $T=(V,E,\mathbf{A})$ be a tree with $N_{\text{leaf}}$ leaves. The tree distance $d_{T}:V\times V\rightarrow\mathbb{R}$ is the sum of weights of the edges on the shortest path between any two nodes on $T$. Let $\Upsilon_{T}(v)$ be the set of nodes in the subtree of the tree $T$ rooted at the node $v\in V$. For any node $u\in V$, there exists a unique parent $v$ s.t. the weight is defined by the tree distance, i.e., $w_{u}=d_{T}(u,v)$ [62]. Given two discrete distributions $\bm{\rho}_{1},\bm{\rho}_{2}\in\mathbb{R}^{N_{\text{leaf}}}$ supported on the tree $T$, the TWD is formally defined by
$\mathrm{TW}(\bm{\rho}_{1},\bm{\rho}_{2},T)=\textstyle\sum_{v\in V}w_{v}\hskip-1.42262pt\left\lvert\sum_{u\in\Upsilon_{T}(v)}\left(\bm{\rho}_{1}(u)-\bm{\rho}_{2}(u)\right)\right\rvert$, (30)
@@
Given a tree $T$ and probability measures supported on the tree $T$, the TWD computing using $T$ is the Wasserstein distance on the tree metric $d_{T}$ [35, 34], i.e.,
@@
However, it is important to note that most of the TWD methods [32, 33, 34, 35, 36] are designed to approximate a Wasserstein distance with an Euclidean ground metric. Specifically, their goal is
@@ A.8 Graph convolutional networks
Graph Convolutional Networks (GCNs) [96] have gained significant attention in graph machine learning, where nodes in a graph are typically assumed in Euclidean spaces. By generalizing convolutional operations to graphs, GCNs effectively capture the dependencies between nodes. This capability enables high accuracy in tasks such as node classification, link prediction, and graph classification, making GCNs useful in applications like social networks [98], molecular structures [124, 125], knowledge graphs [126], recommendation systems [127], and drug discovery [128]. A brief overview of the GCN framework is presented below.
@@
[…]
@@ A.9 Hyperbolic graph convolutional networks
Hyperbolic Graph Convolutional Networks (HGCNs) [21, 45] generalize Graph Convolutional Networks (GCNs) [96] to hyperbolic spaces, using the inductive bias of hyperbolic geometry to better capture hierarchical structures in graph data. By embedding nodes in a negatively curved space, HGCNs provide improved representational capacity for graphs with hierarchies, while preserving the scalability of standard GCNs. This extension requires adapting the steps in GCNs, including feature transformation and neighborhood aggregation, to conform with the geometry of hyperbolic space. HGCNs have shown empirical advantages across a range of tasks, including social network analysis, recommendation systems, and biological network modeling, demonstrating the effectiveness of geometric inductive biases in deep learning on non-Euclidean domains. We briefly review the HGCNs framework below.
@@
Mapping Euclidean node features to hyperbolic representations.
HGCNs [21] begin by mapping input Euclidean node features to a Lorentz model using the exponential map at the origin. Specifically, given a Euclidean feature vector $\mathbf{x}_{i}\in\mathbb{R}^{m}$, its hyperbolic embedding is initialized as
@@
Hyperbolic feature transformation.
To perform feature transformations in hyperbolic space, HGCNs [21] use the logarithmic and exponential maps to move between the manifold and the tangent space at the origin. Let $\mathbf{W}\in\mathbb{R}^{m^{\prime}\times m}$ be a learnable weight matrix. The hyperbolic equivalent of matrix multiplication is defined as
@@
Hyperbolic neighborhood aggregation.
Neighborhood aggregation in HGCNs is performed in the tangent space and leverages hyperbolic attention to account for the hierarchical relationships within the graph. Given hyperbolic embeddings $\mathbf{x}_{i}^{H}$ and $\mathbf{x}_{j}^{H}$, both are first mapped to the tangent space at the origin. The attention weights are computed via an Euclidean Multi-Layer Perceptron ($\mathrm{MLP}$) applied to the concatenation of the tangent vectors
@@
where $[[i]]$ denotes the neighbors of node $i$, and $||$ indicates concatenation. Aggregation is then performed via a hyperbolic weighted average in the tangent space of $\mathbf{x}_{i}^{H}$
@@
Remark A.1.
In the hyperbolic feature transformation step of HGCNs [21], the standard approach involves mapping points between the hyperbolic manifold and the Euclidean tangent space via logarithmic and exponential maps. This allows feature transformations to be performed using Euclidean operations, such as linear transformations followed by bias addition. However, this design only partially uses the advantage of the geometry of hyperbolic space, as the actual transformation takes place in the flat tangent space rather than directly on the curved manifold. To exploit the structure of hyperbolic space, [45] proposed a fully hyperbolic feature transformation. Specifically, they introduce a Lorentz linear transformation that operates directly in the hyperbolic space
@@
Appendix B Extended related works
Manifold learning with Wasserstein distance.
@@ Appendix B Extended related works
Manifold learning is a nonlinear dimensionality reduction approach designed for high-dimensional data that intrinsically lie on a lower-dimensional manifold [58, 51, 114]. Most classical methods derive manifold representations from discrete data samples by constructing a graph, where nodes correspond to data points and edges reflect pairwise similarities. These similarities are often defined using a Gaussian affinity kernel based on the Euclidean norm between the data samples. Recent studies have demonstrated the benefits of using the Wasserstein distance as a more geometry-aware alternative to Euclidean distances in manifold learning frameworks [43, 52], where the Wasserstein distance incorporates the feature relationship. This direction has shown promising results in applications such as matrix organization [129], analysis of neuronal activity patterns [7], and molecular shape space modeling [68]. Building on these ideas, our work explores the integration of the diffusion operator [51] with the TWD to jointly learn hierarchical representations for both samples and features.
Tree-Wasserstein distance.
@@
The Tree-Wasserstein Distance (TWD) offers a computationally efficient alternative to the Wasserstein distance by approximating the ground cost metric with a tree structure. It is designed to compare probability distributions by quantifying the amount of mass that must be transported between them, where the transport is governed by distances along a tree. By approximating the original Euclidean ground metric with a tree metric, TWD enables faster computation while maintaining a close approximation to the true Wasserstein distance. This trade-off between efficiency and fidelity has been validated in a range of studies [32, 33, 34, 35], where TWD has been shown to provide a reliable and scalable surrogate for optimal transport in Euclidean settings. For instance, the Quadtree [32] recursively partitions the ambient space into hypercubes to build random trees, which are then used to calculate the TWD. Flowtree [88] refines this approach by focusing on optimal flow and its cost within the ground metric. Similarly, the Tree-Sliced Wasserstein Distance (TSWD) [35] further improves robustness by averaging TWD values computed over multiple randomly sampled trees, with variants such as TSWD-1, TSWD-5, and TSWD-10 corresponding to the number of sampled trees used. Recent advancements include WQTWD and WCTWD [34], which employ Quadtree or clustering-based tree structures and optimize tree weights to approximate the Wasserstein distance. UltraTree [36] introduces an ultrametric tree by minimizing OT regression cost, aiming to approximate the Wasserstein distance while respecting the original metric space.
@@
Most recently, a tree construction method [37] in TWD literature that differs fundamentally from existing TWD approaches [32, 33, 34, 35]. Whereas conventional TWD methods focus on how TWD can be close to the true Wasserstein distance with the Euclidean ground metric, this method [37] aims to efficiently compute a Wasserstein distance with a ground metric that recovers a latent hierarchical structure underlying hierarchical high-dimensional features. Unlike prior TWD methods that treat tree construction as a heuristic or approximation step to Euclidean distances, this approach [37] treats the tree as a geometric object that reflects intrinsic hierarchical relationships among features. In particular, the constructed tree is not merely a proxy for an ambient Euclidean metric, but rather a data-driven structure whose tree distances approximate geodesic paths on a latent, unobserved hierarchical metric space. This allows the resulting TWD to encode meaningful relationships in settings where the data are governed by hidden hierarchies. In our work, we adopt this tree construction method [37], as it provides a principled and geometry-aware foundation for learning hierarchical representations of both samples and features.
Sliced-Wasserstein distance.
@@
Another widely used approach for efficiently approximating the Wasserstein distance is the Sliced-Wasserstein Distance (SWD) [130, 131]. Instead of solving the high-dimensional optimal transport problem directly, SWD projects the input distributions onto one-dimensional subspaces along random directions, where the Wasserstein distance admits a closed-form solution. Averaging the results over many such slices provides a meaningful measure of similarity while retaining much of the geometric structure inherent in the distributions. This approach is advantageous in high-dimensional settings, where computational efficiency is critical, making it popular in applications such as generative modeling [132], domain adaptation [133], and statistical inference [134].
Co-Manifold learning.
@@
Co-manifold learning aims to simultaneously uncover the manifolds of both samples and features in a data matrix by treating their relationships as mutually informative. That is, the geometry of the features is informed by the samples, and vice versa. This joint modeling approach has been widely explored in tasks such as joint embedding [52, 53, 54, 55] and dimensionality reduction [7, 56, 57], under the assumption that both samples and features lie on smooth, low-dimensional manifolds. By iteratively refining representations across both modes, co-manifold methods aim to capture the manifolds of both samples and features. These approaches typically assume that the manifolds involved are Riemannian and characterized by non-negative curvature [55, 58]. However, such assumptions may be limiting in scenarios where the data exhibit hierarchical organization, which is more naturally modeled in negatively curved spaces such as hyperbolic geometry.
@@
Unsupervised ground metric learning.
Ground metric learning [77] focuses on learning a distance function that defines the cost matrix in the optimal transport problem, thereby capturing meaningful relationships between data elements. The choice of ground metric plays a critical role in downstream tasks [135, 136] such as clustering, transport-based inference, and dimensionality reduction. Many existing approaches rely on prior knowledge of the cost matrix structure or access to label information to guide the learning process [137, 138, 139]. In many cases, however, labeled data is unavailable, making it necessary to develop unsupervised methods that rely solely on the data itself.
Unsupervised ground metric learning seeks to infer a distance metric that captures meaningful pairwise relationships directly from the data, without relying on external labels or prior knowledge. A recent work in this direction is the Wasserstein Singular Vectors (WSV) method [59], which jointly computes Wasserstein distances over samples and features by using the Wasserstein distance matrix in one domain (e.g., samples) as the ground metric for computing distances in the other (e.g., features). However, the method is computationally demanding due to the repeated evaluation of high-dimensional Wasserstein distances. To address this issue, the recent work proposed Tree-WSV [60], which uses tree-based approximations of the Wasserstein distance to reduce computational cost. Both WSV and Tree-WSV adopt an alternating scheme that iteratively updates distances across rows and columns, a strategy that aligns conceptually with our framework.
@@
In contrast, our method is explicitly designed to represent the hierarchical structures for both rows and columns. The key distinction lies in our use of the tree construction method [37] in the TWD: we construct the tree which employs hyperbolic embeddings and diffusion densities to reflect meaningful hierarchical relationships among data points. We then integrate this specific tree-based TWD into the construction of a diffusion operator, whose diffusion densities are used to generate hyperbolic embeddings that encode the joint structure of both samples and features. In doing so, we not only model each hierarchical structure but also capture their mutual influence through an alternating learning process.